OpenStudy (cartoon_horses):

Marsha has a bag that contains 4 green marbles, 8 yellow marbles, and 20 red marbles. If she chooses one marble from the bag, what is the probability that the marble is not red?

OpenStudy (robtobey2):

1-(20/32)

OpenStudy (mariamikayla):

probability=number of favorable cases/number of possible cases the number of favorable cases is obtain by adding the number of green marbles with the number of yellow marbles number of favorable cases : 4+8=12 number of possible casses is the sum of all marbles: 4+8+20=32 probability=12/32= 3/8